40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 3 02:43:34 2013, rana, jenne, Summary, LSC, PRMI: comparison of 1f and 3f signals w/ calibration REFL_signals_CalLinesLinedUp.pdf
    Reply  Mon Oct 7 18:39:29 2013, Jenne, Summary, LSC, PRMI: REFL11 beam realigned 
Message ID: 9218     Entry time: Mon Oct 7 18:39:29 2013     In reply to: 9191
Author: Jenne 
Type: Summary 
Category: LSC 
Subject: PRMI: REFL11 beam realigned 

Quote:

Bonus: notice how we have cleverly used the comb of bounce frequencies around the calibration line to determine that REFL11 is clipping!

 Rana and I noticed last week that it looked like the REFL11 beam was clipping.  This afternoon, I locked the PRMI with REFL 165 I&Q, and checked the REFL 11 path.  The beam looks fine through all of the optics going to the diode, so I just realigned the beam onto the diode using the itty bitty steering mirror.  I have not yet checked the change (hopefully improvement) in the REFL11 spectrum.

ELOG V3.1.3-